The sum of e digits of a two digit number is 9. If the digits are reversed,the new number is 45 less than the original number. Find the original number.
Answer by stanbon(75887)

The sum of the digits of a two digit number is 9. If the digits are reversed,the new number is 45 less than the original number.
Find the original number.
---------------------
Let the number be 10t+u
Reversed you get 10u+t
--------------------------------
Equations:
t + u = 9
10t+u - (10u+t) = 45
----
Rearrange:
t + u = 9
9t-9u = 45
------
Modify;
t + u = 9
t - u = 5
------------------
2t = 14
t = 7
Then u = 2
----
Original Number: 10t+u = 72